OpenStudy (anonymous):

help please! 4^-1=1/4 rewrite as a logarithmic equation

OpenStudy (anonymous):

log (4^-1) = log(1/4) -log (4) = log(1) - log(4) log(1)=0

OpenStudy (anonymous):

well that wasnt hard at all thank ya!

hero (hero):

enjoying yourself?

OpenStudy (anonymous):

so, why you asked if it wasnt hard for you